Merlin has 20 coins in quarters and dimes. The total value of the coins is $3.95. How many coins of each kind does it have? Thank you for you help time!
Answer by josmiceli(19441)

q + d = 20
25q + 10d = 395
25(20-d) + 10d = 395
500 - 25d + 10d = 395
500 - 15d = 395
15d = 500 - 395
15d = 105
d = 7
q = 13
7 dimes and 13 quarters